# Towards Understanding the Benefits of Neural Network Parameterizations in Geophysical Inversions: A Study With Neural Fields

### v1

@@ II-B Proposed NFs-Inv Method
Figure 1: Proposed inversion pipeline. We reparametrize the subsurface model $m$ by the weights of a NN. Considering a 2D DCR inversion, the input of NN f is the 2D spatial coordinates (x, y), and the output is the scalar value of the corresponding physical properties (in this case, conductivity value) on the coordinate. After obtaining physical properties on all coordinates, we have a 2D $m$ that is the predicted subsurface model at the current epoch. This 2D $m$ is then fed into the forward simulation to generate corresponding predicted measurements. After that, the loss function value is computed. During the inverse process, the weights of NN are updated until the loss value is small enough.
A schematic diagram of the proposed methods is shown in Fig. 1. The key component of the proposed method is to re-parameterize the model $m$ by the weights of a multilayer perceptron $f$. In the 2D setting, the input of $f$ is the 2D spatial coordinates of the mesh cell centers (x, z), and the output is the scalar value of the corresponding physical properties that we are inverting for (slowness or log-conductivity) on the coordinate.
The positional encoding is a transform function that maps (x, y) to a higher-dimension space before feeding the input vector into the neural network to overcome spectral bias. Spectral bias is a well-known experimental finding stating that neural networks are biased towards learning lower-frequency variations [47]. This is the reason why a direct mapping from the coordinates space to the physical property space using a neural network has a poor performance in the image representation task—the high-frequency variation has not been learned at the time of convergence, so the reconstructed images are blurred and lacking details [47]. The positional encoding was first proposed in the work of NeRF [8] and then explained by modelling the neural network using a kernel regression with a neural tangent kernel [13]. For context, we will introduce the positional encodings that are relevant to our work. Let $\textbf{x}\in R^{w\times 1}$ denote a vector in the coordinate space, and the output of the encoding function is $\gamma(\textbf{x})\in R^{h\times 1}$. The common positional encoding functions include [13], [8], [10], [48]:
@@
$\begin{split}&\textbf{Gaussian positional encoding}\\ &\gamma(\textbf{x})=[cos(2\pi B\textbf{x}),sin(2\pi B\textbf{x})],\quad\textrm{where B}\in R^{\hat{h}\times\hat{w}}(h=2\hat{h}).\end{split}$ (7)
@@
There are several strategies for defining the trade-off parameter $\beta$. In many cases, a $beta$-cooling strategy is adopted where the inversion starts with a large value for $beta$ and $beta$ is gradually decreased as the inversion progresses in order to fit the data. In the proposed method, we choose to use the same cooling schedule that is employed by the DIP-Inv method [39], where a test-time learning method has been applied to the DCR inversion. The trade-off parameter $\beta$ in the proposed method for the 2D DCR inversion is an exponential decay function shown in equation 10.
@@ II-C Architecture of Neural Networks and Optimization
The number of fully connected layers should be adapted to the complexity of the subsurface model and physics (e.g. it is a function both of the resolution of the forward problem and the complexity of the model). The input dimension of the first hidden layer should depend on the $\gamma$ function we choose. For example, if we choose the basic encoding 5 in 2D, then the input should have dimension 4. For both the cross-hole seismic tomography inversion and DCR inversion examples shown below, the results are obtained using a neural network, $f$, with six hidden layers with the LeakyReLU activation function. The first and last hidden layers have 128 neurons and all other hidden layers have 256 neurons. For the DCR inversions, the activation function for the last layer is a Sigmoid function. For the cross-hole seismic tomography inversions, the activation function for the last layer is a Tanh function to remove the positive constraint provided by the Sigmoid function as a fair comparison to the conventional method. The output in both cases is multiplied by a constant scaler to fit the possible ranges of slowness/conductivity. The details of choosing the $\gamma$ function will be discussed later.
Compared to the conventional method, the proposed method is searching for a solution in a higher-dimension space, so the number of iterations needed to fit the data is larger. The proposed method employs the Adam optimizer, which is a first-order optimization method. The conventional method in the cross-hole seismic tomography inversion employs a gradient descent optimizer, which is also a first-order optimization method. The conventional method in the direct current resistivity inversion employs the Inexact Gaussian Newton method, which is a second-order optimization method. The reason why we do not use a first-order optimization method in the direct current resistivity inversion is that the forward modelling, which is implemented using SimPEG, is much more time-consuming than the cross-hole seismic tomography inversion, so we want to adapt a second-order optimizer that can take fewer iterations to converge. We choose Adam for the proposed method since the adaptive learning rate scheme of Adam is beneficial when updating the weights in the neural networks [49]. A second-order optimization method with an adaptive learning rate has been proposed but is limited by the computational cost of estimating the Hessian when the dimension of the parameter is high [50].
@@ III Trials on the Synthetic Examples
In this section, we will consider two geophysical inversion methods: the cross-hole seismic tomography inversion, a linear problem, and the direct current resistivity (DCR) inversion, a non-linear problem. We present the experiment results on 2D synthetic models to show the benefits of the proposed method. The proposed method can be readily adapted to 3D cases by changing the dimension of the input coordinate from two to three. In the cross-hole seismic tomography inversion, the data have sufficient resolution to recover a model even without a regularization, so we do not use an explicit regularization term in the proposed NF-Inv method. We will compare this result to the conventional method both with and without a L2 smoothness term. In the direct current resistivity inversion, the data has less resolution, so in our proposed method, we use an L1 smallness term. For the conventional method, we will show results that utilize sparse norms on the smallness and smoothness and include sensitivity weighting.
@@ III-A Trials on the Cross-hole Seismic Tomography Inversions
We will consider two models: the first is a simple model of a block in a uniform background, and the second is an elliptical target in a non-homogeneous background. For both cases, the tensor mesh has size $W\times H=64\times 128$, and each cell has a size 1 meter by 1 meter. The two boreholes are separated by 64 meters. The receivers are evenly spaced on one borehole with a separation of 1 meter. The transmitters are evenly spaced on the other borehole with a separation of 1 meter. The physical property model is the slowness (i.e. the reciprocal of velocity).
In the forward simulation, we calculate arrival times (in units of milliseconds), and these are regarded as our geophysical observations. Gaussian noise with a mean of 0 ms and standard deviation of 20 ms has been added to the observations for both cases before inversion. The PDE solver block in Fig. 1 refers to the same forward modelling algorithm.
Fig. 2(a) shows Case 1, a uniform background with a square target. The square target has a velocity of 200 m/s and the background has a velocity of 1000 m/s. In the conventional method, the initial model is chosen to be a uniform background with a value of 1000 m/s. We performed two inversions. Fig. 2(b) shows the case where no explicit regularization was employed during the inverse process, and the recovered model tends to predict structures along the ray path. Fig. 2(c) shows the conventional inversion result with only a smoothness term ($\alpha_{s}=0$, $\alpha_{x}=\alpha_{z}=0.5$, and $p_{x}=p_{z}=2$). By adapting a smoothness term, the unwanted along-ray artifacts have been eliminated, but we can still see some along ray-path artifacts remaining.
Figure 2: The inversion results using the conventional method for Case 1. a) is the true subsurface model. b) is the conventional result without a smoothness term. c) is the conventional result with a L2 smoothness term. The starting/initial model is a uniform half-space for the conventional methods.
@@
Figure 3: The inversion results using the proposed NFs-Inv method for Case 1. a) is the true subsurface model, b) is the starting model, and c) is the recovered subsurface model from the proposed NFs-Inv method without any explicit regularization term.
@@
Figure 4: The inversion results using the conventional method for Case 2. a) is the true subsurface model. b) is the conventional result without any regularization. c) is the conventional result with a smoothness term. The starting/initial model is a uniform half-space for the conventional methods.
Figure 5: The inversion results using the proposed NFs-Inv method for Case 2. a) is the true subsurface model, b) is the starting model, and c) is the recovered subsurface model from the proposed NFs-Inv method without any explicit regularization term.
Similar to Case 1, we performed two inversions. Fig. 4(b) shows the case where no explicit regularization was employed during the inverse process, and the recovered model tends to predict structures along the ray path. Due to the sensitivity and the lack of the positive constraint, the recovered model has unrealistic negative value in some regions. Fig. 4(c) shows the conventional inversion result with only a smoothness term ($\alpha_{s}=0$, $\alpha_{x}=\alpha_{z}=0.5$, and $p_{x}=p_{z}=2$). By adding a smoothness term, the unwanted along-ray artifacts have been eliminated, but the edges of the target are blurred and we can see some along ray-path artifacts remaining.
In the proposed method, the input coordinate space is scaled to [-1, 1] in both the x and z dimensions, and the starting model is randomly initialized [Fig. 5(b)]. To have a better resolution on the inhomogeneous background where the high-frequency variations exist, we choose the Gaussian positional encoding with $B\in R^{128\times 2}$ sampled from a Gaussian distribution with variance 0.5, so the input dimension of $f$ is 256. The impact of the positional encoding will be discussed in Section IV-B. For the Adam optimizer, the learning rate is 0.001. Compared to the recovered models in the conventional method [Fig. 4(b), (c)], the unwanted along-path structures have been eliminated. The heterogeneous background prediction has more details, and the boundaries of the main target are much better resolved [Fig. 5(c)]. Although we can have a better recovery by imposing a positive constraint (e.g. setting the activation function for the last layer to be ReLU), we employ hyperbolic tangent as the activation function in the last layer to have a fair comparison with the conventional method where no positive constraint is imposed. Although there is no positive constraint in both methods, the proposed method performs better than the conventional method in terms of having less prediction of negative values.
@@ III-B Trials on the Direct Current Resistivity Inversions
We again consider two cases, Cases 3 and 4. In both cases, there is a dipping target, but in Case 3, the background is homogeneous, whereas in Case 4, the resistivity of the background increases with depth. For both cases, we choose a dipole-dipole survey geometry with a station separation of 25 meters and a maximum of 24 receivers per transmitter. The survey line is 700 meters long, so the total number of potential difference measurements is 348.
For the forward simulation, we use a mesh that has core cells of 5m $\times$ 5m, with 200 core-mesh cells in the x-direction and 45 cells in the z-direction. We add 7 padding cells that extend by a factor of 1.5 to the sides and bottom of the mesh. In total, the mesh has $W\times H=9000$ active cells. In the proposed method, the input coordinate space is scaled to [-1, 1] in both the x and z dimensions in both cases. The input coordinate space only contains the core domain (active cells), and the padding domain values are set to be the background conductivity and frozen during the inversion process. We use the same positional encoding function (identity $\gamma$ function) and network structure for both cases, and the total number of trainable weights in the neural network in $f$ is 263809, so we over-parameterize the problem.
@@
Figure 6: The inversion results using the proposed method for Case 3. The top row of models (a), are the true models. The models in row b) are from the conventional sparse-norm inversions with sensitivity weighting. The models in row c) are from the DIP-Inv. The models in row d) are from the proposed method.
For the Adam optimizer, the learning rate is 0.001. For the $\beta$-cooling strategy in equation 10, we choose $\tau=800$. The results are shown in [Fig. 6(d)]. Considering another test-time learning method (DIP-Inv [39]) has been developed and tested on the DCR inversion tasks, we also present the inversion result from DIP-Inv for comparison [Fig. 6(c)].
@@
To further test the proposed method in a more complicated scenario, we conduct inversion with the proposed method in a synthetic case with a heterogeneous background [Fig. 7(a)]. The heterogeneous background has a conductivity value that decreases linearly from 0.02 to 0.001 S/m with depth.
@@
The recovered model from the proposed method can be found in Fig. 7(c). For the Adam optimizer, the learning rate is 0.001. Although the boundary of the dike structure cannot be recovered as compactly as in Case 3, the location of the main dike is correctly predicted. The conductivity values for the heterogeneous background are predicted approximately correctly except for the region near the dike structure. Compared to the result from the conventional method, the inversion result from the proposed method has a better depth extent of the dike structure.
Figure 7: a) is the true model for Case 4. b) is the recovered model from the conventional method for Case 4 (with L2 smoothness). c) is the recovered model from the NFs-Inv method for Case 4 (without any explicit regularization).
@@ IV-A Benefits of Changing the Searching Space
In Section III-A, the results demonstrate that when searching in the space of the NN-weights for a solution, the inversion results are less influenced by sensitivity artifacts. This is especially noticeable when comparing the conventional and NFs-Inv results when no explicit regularization is used [Fig. 8]. Unlike the seismic inversions, the conventional DCR inversion results are not very meaningful if no regularization is included [Fig. 8(b)]. Namely, searching for the solution in the original mesh space is largely affected by the sensitivity, so the results have unrealistic artifacts near the electrodes [Fig. 8(b)] and therefore the conventional approach cannot correctly predict the structures at depth, such as the dike in Case 3. However, even if we do not apply any regularization to the proposed method (i.e. $\beta=0$), we are able to recover the main features, for example, the dike and layered structures in Case 3 [8(c)]. Considering that the only regularization is from the implicit bias of the proposed method, it is fair to conclude that the neural field methods have a useful implicit regularization effect on the geophysical inverse problem. The challenge of overcoming artifacts due to the sensitivities of a given survey is common across many geophysical experiments (e.g., it is common to see structures cluster around ground electromagnetic stations). This also indicates the potential application of the NFs-Inv method to other geophysical data types.
Figure 8: a) is the true model for Case 3. b) is the recovered model from the conventional method without any explicit regularization for Case 3. c) is the recovered model from the proposed NFs-Inv method without any explicit regularization for Case 3.
IV-B Apply Positional Encoding to Recovery the Higher-frequency Variances

### v2

@@ II-B Proposed NFs-Inv Method
Figure 1: Proposed inversion pipeline. We reparametrize the subsurface model $m$ by the weights of a NN. Considering a 2D DCR inversion, the input of NN f is the 2D spatial coordinates (x, y), and the output is the scalar value of the corresponding physical property (in this case, the conductivity value) at the coordinate. After obtaining physical properties on all coordinates, we have a 2D $m$ that is the predicted subsurface model at the current epoch. This 2D $m$ is then fed into the forward simulation to generate corresponding predicted measurements. After that, the loss function value is computed. During the inverse process, the weights of the NN are updated until the loss value is small enough.
A schematic diagram of the proposed methods is shown in Fig. 1. The key component of the proposed method is to reparameterize the model $m$ by the weights of a multilayer perceptron $f$. In the 2D setting, the input of $f$ is the 2D spatial coordinates of the mesh cell centers (x, z), and the output is the scalar value of the corresponding physical properties for which we invert (slowness or log-conductivity) on the coordinate.
Positional encoding is a transform function that maps (x, y) to a higher-dimensional space before feeding the input vector into the neural network to overcome spectral bias. Spectral bias is a well-known experimental finding stating that neural networks are biased towards learning lower-frequency variations [55]. This is the reason why a direct mapping from the coordinate space to the physical property space using a neural network has poor performance in the image representation task—the high-frequency variation has not been learned at the time of convergence, so the reconstructed images are blurred and lack details [55]. Positional encoding was first proposed in the work of NeRF [11] and then explained using the neural tangent kernel theory [17]. For context, we will introduce the positional encodings that are relevant to our work. Let $\textbf{x}\in R^{w\times 1}$ denote a vector in the coordinate space, and the output of the encoding function is $\gamma(\textbf{x})\in R^{h\times 1}$. The common positional encoding functions include [11], [13], [17], [56]:
@@
$\begin{split}&\textbf{Gaussian positional encoding}\\ &\gamma(\textbf{x})=[cos(2\pi B\textbf{x}),sin(2\pi B\textbf{x})],\\ &\textrm{where B}\in R^{\hat{h}\times\hat{w}}(h=2\hat{h})\textrm{ is sampled from a Gaussian distribution}.\end{split}$ (7)
@@
There are several strategies for defining the trade-off parameter $\beta$. In many cases, a $\beta$-cooling strategy is adopted where the inversion starts with a large value for $\beta$ and $\beta$ is gradually decreased as the inversion progresses in order to fit the data. In the proposed method, we choose to use the same cooling schedule that is employed by the DIP-Inv method [47], where a test-time learning method has been applied to the DCR inversion. The trade-off parameter $\beta$ in the proposed method for the 2D DCR inversion is an exponential decay function shown in equation 10.
@@ II-C Architecture of Neural Networks and Optimization
The number of fully connected layers should be adapted to the complexity of the subsurface model and physics (e.g. it is a function of both the resolution of the forward problem and the complexity of the subsurface model). The input dimension of the first hidden layer should depend on the $\gamma$ function we choose. For example, if we choose the basic encoding 5 in 2D, then the input should have dimension 4. For both the cross-hole seismic tomography inversion and DCR inversion examples shown below, the results are obtained using a neural network, $f$, with six hidden layers with the LeakyReLU activation function. The first and last hidden layers have 128 neurons, and all other hidden layers have 256 neurons. For the DCR inversions, where we search the solution in the log-conductivity space, the activation function for the last layer is a Sigmoid function. For the cross-hole seismic tomography inversions, the activation function for the last layer is a Tanh function. The output in both cases is multiplied by a constant scalar to fit the possible ranges of slowness/conductivity. The details of choosing the $\gamma$ function will be discussed in Section III.
Compared to the conventional method, the proposed method searches for a solution in a higher-dimensional space, so the number of iterations needed to fit the data is larger. The proposed method employs the Adam optimizer, which is a first-order optimization method. The conventional method in the cross-hole seismic tomography inversion employs a gradient descent optimizer, which is also a first-order optimization method. The conventional method in the direct current resistivity inversion employs the Inexact Gaussian Newton method, which is a second-order optimization method. The reason why we do not use a first-order optimization method in the direct current resistivity inversion is that the forward modelling, which is implemented using SimPEG, is much more time-consuming than the cross-hole seismic tomography inversion, so we want to adapt a second-order optimizer that can take fewer iterations to converge. We choose Adam for the proposed method since the adaptive learning rate scheme of Adam is beneficial when updating the weights in the neural networks [57]. A second-order optimization method with an adaptive learning rate has been proposed but is limited by the computational cost of estimating the Hessian when the dimension of the parameter is high [58].
@@ III Trials on the Synthetic Examples
In this section, we will consider two geophysical inversion methods: the cross-hole seismic tomography inversion, a linear problem, and the direct current resistivity (DCR) inversion, a non-linear problem. We present the experiment results on 2D synthetic models to show the benefits of the proposed method. The proposed method can be readily adapted to 3D cases by changing the dimension of the input coordinate from two to three. In the cross-hole seismic tomography inversion, the data has sufficient resolution to recover a model even without a regularization, so we do not use an explicit regularization term in the proposed NF-Inv method. We will compare this result to the conventional method, both with and without an L2 smoothness term. In the direct current resistivity inversion, the data has less resolution, so in our proposed method, we use an L1 smallness term. For the conventional method, we will show results that utilize sparse norms on the smallness and smoothness and include a sensitivity weighting in the objective function.
@@ III-A Trials on the Cross-hole Seismic Tomography Inversions
We will consider two models: the first is a simple model of a block in a uniform background, and the second is an elliptical target in a non-homogeneous background. For both cases, the tensor mesh has size $W\times H=64\times 128$, and each cell has a size of 1 meter by 1 meter. The two boreholes are separated by 64 meters. The receivers are evenly spaced on one borehole with a separation of 1 meter. The transmitters are evenly spaced on the other borehole with a separation of 1 meter.
In the forward simulation, we calculate arrival times (in units of milliseconds), and these are regarded as our geophysical observations. Gaussian noise with a mean of 0 ms and a standard deviation of 20 ms has been added to the observations for both cases before inversion. The PDE solver block in Fig. 1 refers to the linear forward modelling algorithm.
Fig. 2(a) shows Case 1, a uniform background with a square target. The square target has a velocity of 200 m/s, and the background has a velocity of 1000 m/s. In the conventional method, the initial model is chosen to be a uniform background with a value of 1000 m/s. We perform two conventional inversions. Fig. 2(b) shows the case where no explicit regularization was employed during the inverse process, and the recovered model tends to predict structures along the ray path. Fig. 2(c) shows the conventional inversion result with only a smoothness term ($\alpha_{s}=0$, $\alpha_{x}=\alpha_{z}=0.5$, and $p_{x}=p_{z}=2$). By adapting a smoothness term, the unwanted along-ray artifacts have been eliminated, but we can still see some along-ray-path artifacts remaining.
Figure 2: The inversion results using the conventional method for Case 1. (a) is the true subsurface model. (b) is the conventional result without a smoothness term. (c) is the conventional result with the L2 smoothness term. The starting/initial model is a uniform half-space for the conventional methods.
@@
Figure 3: The inversion results using the proposed NFs-Inv method for Case 1. (a) is the true subsurface model. (b) is the starting model. (c) is the recovered subsurface model from the proposed NFs-Inv method without any explicit regularization term.
@@
Figure 4: The inversion results using the conventional method for Case 2. (a) is the true subsurface model. (b) is the conventional result without any regularization. (c) is the conventional result with the L2 smoothness term. The starting/initial model is a uniform half-space for the conventional methods.
Figure 5: The inversion results using the proposed NFs-Inv method for Case 2. (a) is the true subsurface model, (b) is the starting model, and (c) is the recovered subsurface model from the proposed NFs-Inv method without any explicit regularization term.
Similar to Case 1, we performed two inversions. Fig. 4(b) shows the case where no explicit regularization was employed during the inverse process, and the recovered model tends to predict structures along the ray path. Due to the sensitivity and the lack of a positivity constraint, the recovered model has unrealistic negative values in some regions. Fig. 4(c) shows the conventional inversion result with only a smoothness term ($\alpha_{s}=0$, $\alpha_{x}=\alpha_{z}=0.5$, and $p_{x}=p_{z}=2$). By adding a smoothness term, the unwanted along-ray artifacts have been eliminated, but the edges of the target are blurred, and we can see some along-ray-path artifacts remaining.
In the proposed method, the input coordinate space is scaled to [-1, 1] in both the x and z dimensions, and the starting model is randomly initialized [Fig. 5(b)]. To have a better resolution on the inhomogeneous background where the high-frequency variations exist, we choose the Gaussian positional encoding with $B\in R^{128\times 2}$ sampled from a Gaussian distribution with a standard deviation 0.5, so the input dimension of $f$ is 256. The impact of the positional encoding will be discussed in Section IV-B. For the Adam optimizer, the learning rate is 0.001. Compared to the recovered models in the conventional method [Fig. 4(b), (c)], the NFs-Inv does not have any unwanted along-path structures. The heterogeneous background prediction has more details, and the boundaries of the main target are much better resolved [Fig. 5(c)]. Although we can have a better recovery by imposing a positive constraint (e.g. setting the activation function for the last layer to be ReLU), we employ the hyperbolic tangent as the activation function in the last layer to have a fair comparison with the conventional method where no positive constraint is imposed. Although there is no positive constraint in both methods, the proposed method performs better than the conventional method in terms of having fewer predictions of negative values.
@@ III-B Trials on the Direct Current Resistivity Inversions
We again consider two cases (Cases 3 and 4). In both cases, there is a dipping target, but in Case 3, the background is homogeneous, whereas in Case 4, the resistivity of the background increases with depth. For both cases, we choose a dipole-dipole survey geometry with a station separation of 25 meters and a maximum of 24 receivers per transmitter. The survey line is 700 meters long, so the total number of potential difference measurements is 348.
For the forward simulation, we use a mesh that has core cells of 5m $\times$ 5m, with 200 core-mesh cells in the x-direction and 45 cells in the z-direction. We add 7 padding cells that extend by a factor of 1.5 to the sides and bottom of the mesh. In total, the mesh has $W\times H=9000$ active cells. In the proposed method, the input coordinate space is scaled to [-1, 1] in both the x and z dimensions in both cases. The input coordinate space only contains the core domain (active cells), and the padding domain values are set to be the background conductivity and frozen during the inversion process. We use the same positional encoding function (identity $\gamma$ function) and network structure for both cases, and the total number of trainable weights in the neural network in $f$ is 263809.
@@
Figure 6: The inversion results using the proposed method for Case 3. The top row (a) is the true model. The models in row (b) are from the conventional sparse-norm inversions with sensitivity weighting. The models in row (c) are from the DIP-Inv. The models in row (d) are from the proposed method.
For the Adam optimizer, the learning rate is 0.001. For the $\beta$-cooling strategy in equation 10, we choose $\tau=800$. The results are shown in Fig. 6(d). Considering another test-time learning method (DIP-Inv [47]) has been developed and tested on the DCR inversion tasks, we also present the inversion result from DIP-Inv for comparison [Fig. 6(c)].
@@
To further test the proposed method in a more complicated scenario, we conduct an inversion with the proposed method in a synthetic case with a heterogeneous background [Fig. 7(a)]. The heterogeneous background has a conductivity value that decreases linearly from 0.02 to 0.001 S/m with depth.
@@
The recovered model from the proposed method can be found in Fig. 7(c). For the Adam optimizer, the learning rate is 0.001. Although the boundary of the dike structure cannot be recovered as compactly as in Case 3, the location of the main dike is correctly predicted. The conductivity values for the heterogeneous background are predicted approximately correctly, except for the region near the dike structure. Compared to the result from the conventional method, the inversion result from the proposed method has a better depth extent of the dike structure.
Figure 7: (a) is the true model for Case 4. (b) is the recovered model from the conventional method for Case 4 (with L2 smoothness). (c) is the recovered model from the NFs-Inv method for Case 4 (without any explicit regularization).
@@ IV-A Benefits of Changing the Searching Space
In Section III-A, the results demonstrate that when searching in the space of the NN-weights for a solution, the inversion results are less influenced by sensitivity artifacts. This is especially noticeable when comparing the conventional and NFs-Inv results when no explicit regularization is used [Fig. 8]. Unlike the seismic inversions, the conventional DCR inversion results are not very meaningful if no regularization is included [Fig. 8(b)]. Namely, searching for the solution in the original mesh space is largely affected by the sensitivity, so the results have unrealistic artifacts near the electrodes [Fig. 8(b)] and therefore the conventional approach cannot correctly predict the structures at depth, such as the dike in Case 3. However, even if we do not apply any regularization to the proposed method (i.e. $\beta=0$), we are able to recover the main features, for example, the dike and layered structures in Case 3 [8(c)]. Considering that the only regularization is from the implicit bias of the proposed method, it is fair to conclude that the neural fields methods have a useful implicit regularization effect on the geophysical inverse problem. The challenge of overcoming artifacts due to the sensitivities of a given survey is common across many geophysical experiments (e.g., it is common to see structures cluster around ground electromagnetic stations). This also indicates the potential application of the NFs-Inv method to other geophysical data types.
Figure 8: (a) is the true model for Case 3. (b) is the recovered model from the conventional method without any explicit regularization for Case 3. (c) is the recovered model from the proposed NFs-Inv method without any explicit regularization for Case 3.
IV-B Apply Positional Encoding to Recover the Higher-frequency Variation